# Topological consequences of null-geodesic refocusing and applications to $Z^{x}$ manifolds

### v1

@@ 1 Introduction
Let $(M,h)$ be a complete Riemannian manifold, $x\in M$ and $l>0$. We say that $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold if for all geodesics $\alpha$ in $(M,h)$ with $\alpha(0)=x$ and $g(\alpha^{\prime}(0),\alpha^{\prime}(0))=1$ we have $\alpha(l)=y$. If $x=y$ we call $(M,h)$ a $Y^{x}_{l}$ manifold. Note also that every $Y^{(x,y)}_{l}$ is a $Y^{x}_{2l}$ manifold (see [12]).
@@
Let $(M,h)$ be a complete Riemannian manifold, $x\in M$ We say that $(M,h)$ is a $Z^{(x,y)}$ manifold if for all geodesics $\alpha$ in $(M,h)$ with $\alpha(0)=x$ there is some $t>0$ with $\alpha(t)=y$. If $x=y$ we call $(M,h)$ a $Z^{x}$ manifold. In contrast to $Y^{(x,y)}_{l}$ manifolds it is not known whether every $Z^{(x,y)}$ manifold is a $Z^{x}$ manifold.
@@
1.4 Definition (observer refocusing spacetime).
Let $(X,g)$ be a spacetime, $p\in X$ and $\gamma$ some timelike curve in $X$. We say that $(X,g)$ is observer-refocusing if for all future-oriented null-geodesics $\alpha$ in $(X,g)$ with $\alpha(0)=x$ there is some $t>0$ with $\alpha(t)\in\mathrm{im}(\gamma)$.
@@
For a spacetime $(X,g)$ a smooth temporal function $\mathcal{T}:X\to\mathbb{R}$ is a smooth function with everywhere future-pointing timelike gradient. In [3] [4] Bernal and Sanchez prove that for every globally hyperbolic spacetime $X$, there are smooth, surjective temporal functions $\mathcal{T}:X\to\mathbb{R}$ such that $S_{t}\coloneqq\mathcal{T}^{-1}(t)\subset X$ is a smooth spacelike Cauchy surface for each $t\in\mathbb{R}$. Smooth temporal functions are strictly increasing along every future-directed causal curve.
@@ 2 Main Results
3 Related results about $Y^{x}_{l}$ manifolds and weaker notions of refocusing spacetimes
We first review results about the topology of $Y^{x}_{l}$ manifolds and refocusing spacetimes. $Y^{x}_{l}$ manifolds, as well as $Z^{x}$ manifolds and a whole family of related notions, are discussed in [8, Definitions 7.7]. The following result is due to Bott [9] and Samelson [26].
@@ 3 Related results about $Y^{x}_{l}$ manifolds and weaker notions of refocusing spacetimes
By recent contact Bott-Samelson Theorem results of Frauenfelder, Labrousse and Schlenk [13], the conclusion of the Bérard-Bergery theorem can be strengthened to be exactly the same as the one in Bott-Samelson Theorem. Note that is it not know if every $Y^{x}_{l}$ manifold admits a metric in which all geodesics at $x$ are periodic.
@@
Low proved that globally hyperbolic spacetimes must have compact Cauchy surface. In [12, Theorem 11.5] Chernov and Rudyak prove that globally hyperbolic refocusing spacetimes of dimension at least $3$ have a globally hyperbolic semi-Riemannian universal cover which is also refocusing. They use this to prove the following result.
@@
Just as it is unknown whether there are $Z^{x}$ manifolds that are not $Y^{x}_{l}$ manifolds [8, Question 7.70], it is unclear whether there are $\tilde{Y}^{x}$ manifolds that are not $Y^{x}_{l}$ manifolds, whether there are refocusing spacetimes that are not strongly refocusing, and whether there are observer-refocusing spacetimes that are not strongly refocusing.
@@
Every $Z^{x}$ manifold is a $\tilde{Z}^{x}$ manifold but the converse is false. For example, any $n$-torus $T^{n}$ with standard metric will be a $\tilde{Z}^{x}$ manifold, but only $T^{1}$ is a $Z^{x}$ manifold. Similarly there are not observer-refocusing. We will show in Proposition 4.29 that globally hyperbolic very weakly refocusing spacetimes have compact Cauchy surface. In fact every proof of compactness of Cauchy surfaces in this paper will ultimately rely on this face.
@@ 4 Observer-Refocusing Spacetimes
Since the result is entirely local, we can assume $M=\mathbb{R}^{k}$ and $x=0$. Without loss of generality, assume that $x_{n}\to x$ and that $x_{n}\neq x$ for all $n$, else take a subsequence. Further, without loss of generality, assume that $\lim_{n\to\infty}\frac{x_{n}}{||x_{n}||}=v$ for some unit vector $v$ and that $||x_{n+1}||<\frac{1}{2^{n}}||x_{n}||$ for all $n$, else we again take a subsequence. Then the sequence $(w_{n})_{n\in\mathbb{N}}$ defined through $w_{n}\coloneqq\frac{x_{n}-x_{n+1}}{||x_{n}-x_{n+1}||}$ converges to $v$. We define the curve $\gamma$ as follows. We connect $x_{n+1}$ to $x_{n}$ by a a $\mathcal{C}^{1}$ curve $\gamma_{n}:[0,a_{n}]$, parametrized by arc-length, satisfying $\gamma_{n}(0)=x_{n+1}$, $\gamma_{n}(a_{n})=x_{n}$, $\gamma_{n}^{\prime}(0)=w_{n+1}$ and $\gamma_{n}^{\prime}(a_{n})=w_{n}$. Since the sequence $(w_{n})_{n\in\mathbb{N}}$ converges to $v$, the interpolations can be done in a way (for example by interpolating with cubic splines) such that that $\lim_{n\to\infty}\sup_{t\in[0.a_{n}]}||\gamma_{n}^{\prime}(t)-v||=0$ and that $\sum_{n\in\mathbb{N}}a_{n}<\infty$. We then define $\gamma$ as the infinite concatenation
@@
Let $(X,g)$ be a Lorentzian manifold. $h:[0,1]\to T_{p}X$ a curve such that $h(t)$ is lightlike for each $t\in[0,1]$. Then $h^{\prime}(0)$ is not timelike. Hence $d\exp_{0}(h^{\prime}(0))=\frac{d}{dt}|_{t=0}\exp(h(t))$ is not timelike.
@@
To show that the Cauchy surface of $(X,g)$ has finite fundamental group, we proceed as follows. Let $\rho:(\tilde{X},\tilde{g})\to(X,g)$ be the semi-Riemannian universal cover of $X$. Let $\tilde{p}$ be any point in $\rho^{-1}(p)$ and let $\tilde{L}$ be the lightsphere at $p$ given by $\tilde{L}=(d\rho)^{-1}(L)$. Note that the set $\rho^{-1}(\{q_{i}\}_{i\in I}\})$ is countable. Let $\tilde{v}\in\tilde{L}$ and $v=d\rho(\tilde{v})\in L$. Then there is some $t\in(0,T]$ and $i\in I$ with $\alpha_{v}(t)=q_{i}$. Lifting $\alpha_{v}$ to $\tilde{X}$, we get that $\alpha_{\tilde{v}}(t)\in\rho^{-1}(q_{i})\in\rho^{-1}(\{q_{i}\}_{i\in I}\})$. Note that $\rho^{-1}(\{q_{i}\}_{i\in I}\})$ is countable. Through application of the first part of this theorem, we can conclude that $\tilde{X}$ has compact Cauchy surface as well. Now pick a Cauchy surface $S\subset X$ and let $\tilde{S}=\rho^{-1}(S)$. Then $\tilde{S}$ is a Cauchy surface of $\tilde{X}$ (see the proof of [12, Theorem 11.5]) and hence compact. Further, $\tilde{S}$ is simply connected (as $\tilde{X}$ is simply connected and diffeomorphic to $\tilde{S}\times\mathbb{R}$), and $\rho:\tilde{S}\to S$ is a covering map. Pick any $x\in\tilde{S}$, then
@@
It is unknown whether there are $Z^{x}$ manifolds where the time-of-return of unit-speed geodesics is not uniformly bounded [8, Question 7.71]. Similarly it is unknown whether there are any globally hyperbolic observer-refocusing spacetimes where the timelike curve cannot be taken to be compact.
@@ 5 The Analytic Case
Let $(X,g)$ be a globally hyperbolic spacetime with $g$ an analytic metric. Then there exists an analytic, surjective temporal function $\mathcal{T}:X\to\mathbb{R}$ such that every level set of $\mathcal{T}$ is a spacelike Cauchy hypersurface.
@@ 6 Analogous conjecture in the contact case
The results of Chernov and Nemirovski [11], the results of Frauenfelder, Labrousse and Schlenk [13, Theorem 1.11], and our results about observer-refocusing spacetimes and Riemannian $Z^{x}$ manifolds in this paper, suggest the following conjectures.
@@
Let $M$ be manifold of dimension at least $2$, let $(S^{\ast}M,\xi)$ be the spherical cotangent bundle on $M$ equipped with the standard contact structure. Let $\alpha$ be any contact form on $(S^{\ast}M,\xi)$, i.e. $\xi=\ker(\alpha)$, and let $\varphi_{\alpha}^{t}$ be the Reeb flow of $\alpha$. Assume that for all $v\in S^{\ast}_{x}M$ there is some $t>0$ with $\varphi_{\alpha}^{t}(v)\in S^{\ast}_{x}M$. Then $M$ is compact, the fundamental group of $M$ is finite, and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
@@
Let $M$, $\xi$, $\alpha$ and $\varphi_{\alpha}^{t}$ be as in Conjecture 6.46. Let $x\in M$. Assume there is some $T>0$ such that for all $v\in S^{\ast}_{x}M$ there is some $t\in(0,T]$ with $\varphi_{\alpha}^{t}(v)\in S^{\ast}_{x}M$. Then $M$ is compact, the fundamental group of $M$ is finite and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
@@
This contact-geometric version generalizes the Riemannian $Z^{x}$ version as follows: Let $(M,g)$ be a Riemannian manifold, let $\psi:TM\to T^{\ast}M$ be the isomorphism defined through $v\mapsto g(v,\cdot)$, then $\psi$ induces a bundle-isomorphism (which we will also denote by $\psi$) between the unit-tangent bundle $U_{g}M$ and the spherical cotangent bundle $S^{\ast}M$. We can define a $1$-form $\alpha$ on $S^{\ast}M$ via $\alpha_{v}(w)=g(\psi^{-1}(v),d\pi(w))$ where $\pi:S^{\ast}M\to M$ is the natural projection. The Reeb flow of $\alpha$ on $S^{\ast}M$ corresponds to the co-geodesic flow of $g$ on $S^{\ast}M$. Conjecture 6.46 generalizes the conjecture that $Z^{x}$ manifolds of dimension at least $2$ have finite fundamental group. We have proven the compactness and the fundamental group aspect of Conjecture 6.47 whenever the $1$-forms $\alpha$ comes from a Riemannian metric. The conjecture states that this is really a fact about Reeb flows on $(S^{\ast}M,\xi)$ and not just about geodesics on $M$. One may also ask other questions, for example if the set $R_{x}^{\alpha}\coloneqq\{t>0\mid\exists v\in S^{\ast}_{x}M\text{ such that }\varphi_{\alpha}^{t}(x)\in S^{\ast}_{x}M\}$ is discrete for all contact forms $\alpha$ on $(S^{\ast}M,\xi)$. If it is always discrete, then Conjecture 6.46 holds for analytic $1$-forms $\alpha$ by the analyticity of the Reeb flow and by the analyticity of $\pi:S^{\ast}M\to M$ because arguments as in Theorem 5.44 can reduce the problem to the case already solved by Frauenfelder, Labrousse and Schlenk.

### v2

@@ 1 Introduction
Let $(M,h)$ be a complete Riemannian manifold, $x,y\in M$ and $l>0$. We say that $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold if for all geodesics $\alpha$ in $(M,h)$ with $\alpha(0)=x$ and $g(\alpha^{\prime}(0),\alpha^{\prime}(0))=1$ we have $\alpha(l)=y$. If $x=y$ we call $(M,h)$ a $Y^{x}_{l}$ manifold. Note also that every $Y^{(x,y)}_{l}$ is a $Y^{x}_{2l}$ manifold (see [15]).
@@
Let $(M,h)$ be a complete Riemannian manifold, $x,y\in M$ We say that $(M,h)$ is a $Z^{(x,y)}$ manifold if for all geodesics $\alpha$ in $(M,h)$ with $\alpha(0)=x$ there is some $t>0$ with $\alpha(t)=y$. If $x=y$ we call $(M,h)$ a $Z^{x}$ manifold. In contrast to $Y^{(x,y)}_{l}$ manifolds it is not known whether every $Z^{(x,y)}$ manifold is a $Z^{x}$ manifold.
@@
1.4 Definition (observer-refocusing spacetime).
Let $(X,g)$ be a spacetime, $p\in X$ and $\gamma$ some timelike curve in $X$. We say that $(X,g)$ is observer-refocusing if for all future-oriented null-geodesics $\alpha$ in $(X,g)$ with $\alpha(0)=p$ there is some $t>0$ with $\alpha(t)\in\mathrm{im}(\gamma)$.
@@
For a spacetime $(X,g)$ a smooth temporal function $\mathcal{T}:X\to\mathbb{R}$ is a smooth function with everywhere past-pointing timelike gradient. In [4], [5] Bernal and Sanchez prove that for every globally hyperbolic spacetime $X$, there are smooth, surjective temporal functions $\mathcal{T}:X\to\mathbb{R}$ such that $S_{t}\coloneqq\mathcal{T}^{-1}(t)\subset X$ is a smooth spacelike Cauchy surface for each $t\in\mathbb{R}$. Smooth temporal functions are strictly increasing along every future-directed causal curve.
@@ 2 Main Results
3 Related results about $Y^{x}_{l}$ manifolds and refocusing spacetimes
We first review results about the topology of $Y^{x}_{l}$ manifolds and refocusing spacetimes. $Y^{x}_{l}$ manifolds, as well as $Z^{x}$ manifolds and a whole family of related notions, are discussed in [9, Definitions 7.7]. The following result is due to Bott [10] and Samelson [29].
@@ 3 Related results about $Y^{x}_{l}$ manifolds and refocusing spacetimes
By recent contact Bott-Samelson Theorem results of Frauenfelder, Labrousse and Schlenk [16], the conclusion of the Bérard-Bergery theorem can be strengthened to be exactly the same as the one in Bott-Samelson Theorem. Note that is it not known if every $Y^{x}_{l}$ manifold admits a metric in which all geodesics at $x$ are periodic.
@@
Low proved that globally hyperbolic spacetimes must have compact Cauchy surface. In [15, Theorem 11.5] Chernov and Rudyak prove that globally hyperbolic refocusing spacetimes of dimension at least $3$ have a globally hyperbolic semi-Riemannian universal cover which is also refocusing. They use this to prove the following result.
@@
Just as it is unknown whether there are $Z^{x}$ manifolds that are not $Y^{x}_{l}$ manifolds [9, Question 7.70], it is unclear whether there are $\tilde{Y}^{x}$ manifolds that are not $Y^{x}_{l}$ manifolds, whether there are refocusing spacetimes that are not strongly refocusing, and whether there are observer-refocusing spacetimes that are not strongly refocusing.
@@
Every $Z^{x}$ manifold is a $\tilde{Z}^{x}$ manifold but the converse is false. For example, any $n$-torus $T^{n}$ with standard metric will be a $\tilde{Z}^{x}$ manifold, but only $T^{1}$ is a $Z^{x}$ manifold. Similarly there are not observer-refocusing. We will show in Proposition 4.32 that globally hyperbolic very weakly refocusing spacetimes have compact Cauchy surface. In fact every proof of compactness of Cauchy surfaces in this paper will ultimately rely on this fact.
@@
In [22] Low showed that the space of future-directed, unparameterized, maximal null-geodesics $\mathcal{N}$ of a strongly causal spacetime $(X,g)$ is a smooth manifold with a natural contact structure. For globally hyperbolic manifolds, there is a natural contactomorphism from $\mathcal{N}$ to $S^{\ast}M$ for any spacelike Cauchy surface $M\subset X$ given by the following map: To each future-directed null-geodesic $\alpha$ in $X$ there is a unique point $p\in M$ and $t\in\mathbb{R}$ with $\alpha(t)=p$. Let $\pi:X=M\times\mathbb{R}\to M$ be the natural projection. Consider the vector $v\in T_{p}M$ defined through $v=d\pi(\alpha^{\prime}(t))$. Orientation preserving affine reparameterization will change $v$ by multiplication with a positive scalar. Let $[\alpha]$ denote the equivalence class of null-geodesics up to reparametrization and let $[v]$ denote the equivalence class of vectors in $TM$ up to multiplication by a positive scalar. Then the map $\mathcal{N}\to SM$, $[\alpha]\mapsto[v]$ is well-defined. It is easy to check that this map is a diffeomorphism. Since $g$ restricts to a Riemannian metric on $M$, we can identify $SM$ with $S^{\ast}M$ through $g|_{M}$. This gives us the contactomorphism $\mathcal{N}\to S^{\ast}M$. For each $p\in X$ the set $\mathrm{sky}_{p}\coloneqq\{[\alpha]\in\mathcal{N}\mid\alpha\text{ is a null-geodesic through }p\}$ is called the sky of $p$. Each sky is a Legendrian submanifold of $\mathcal{N}$, and if $M$ is any Cauchy surface through $p$ then $\mathrm{sky}_{p}$ corresponds to a fiber of $S^{\ast}M$ when identifying $\mathcal{N}\simeq S^{\ast}M$.
3.26 Proposition.
Let $X$ be a globally hyperbolic spacetime which splits as $X=M\times\mathbb{R}$ with metric $g=h-Adt^{2}$ where $A$ is some smooth, positive function and where $h$ is a symmetric $2$-tensor with $h(\partial_{t},\cdot)=0$ and such that $h$ restricts to a Riemannian metric on $M\times\{t\}$ for each $t\in\mathbb{R}$. For $t\in\mathbb{R}$ let $h_{t}$ denote the metric on $M\simeq M\times\{t\}$ given by $h_{t}=h|_{M\times\{t\}}$ and let $\alpha_{t}$ be the contact form on $(S^{\ast}M,\xi)$ induced by $h_{t}$. The co-orientations induced by each $\alpha_{t}$ agree. For each $t\in\mathbb{R}$, let $\phi_{t}:\mathcal{N}\to S^{\ast}(M\times\{t\})\simeq S^{\ast}M$ be the natural contactomorphism. Let $p\in X$. Then $\phi_{t}(\mathrm{sky}_{p})$ is a positive Legendrian isotopy in $S^{\ast}M$ (with respect to the co-orientation of any of the contact forms $\alpha_{t}$).
Proof.
That $(\phi_{t}(\mathrm{sky}_{p}))_{t\in\mathbb{R}}$ is a Legendrian isotopy follows from the fact that each $\phi_{t}$ is a contactomorphism. It is left to show that the isotopy is positive. Let $[\beta]\in\mathrm{sky}_{p}$. We can parametrize $\beta$ as a pre-geodesic so that $\beta(t)=(x(t),t)$. Note that $\pi(\phi_{t}([\beta]))=x(t)$. Let $\psi:U_{h_{t}}M\to S^{\ast}M$ be the natural map from the unit tangent bundle to the spherical cotangent bundle. By definition of $\phi_{t}$, there is some $\lambda>0$ with $\psi^{-1}(\phi_{t}([\beta]))=\lambda x^{\prime}(t)$. Therefore
$\alpha_{t}\left(\frac{d}{dt}\phi_{t}([\beta])\right)$ $=h_{t}\left(\psi^{-1}(\phi_{t}([\beta])),d\pi\left(\frac{d}{dt}\phi_{t}([\beta])\right)\right)$
$=h_{t}\left(\lambda x^{\prime}(t),x^{\prime}(t)\right)>0.$
We can now prove a Bott-Samelson type result for the Cauchy surfaces of globally hyperbolic strongly refocusing spacetimes.
3.27 Proposition.
Let $X$ be a globally hyperbolic, strongly refocusing spacetimes of dimension at least $3$. Then the Cauchy surface $M$ of $X$ is compact with finite fundamental group, and its universal cover has the integral cohomology ring of a CROSS.
Proof.
The construction of Proposition 3.26 shows that there is a positive Legendrian isotopy from a fiber $S_{x}^{\ast}M$ to a fiber $S_{y}^{\ast}M$ of $S^{\ast}M$. Since $S^{\ast}M$ is symmetric with respect to choice of fibers, there also is a positive Legendrian isotopy from $S_{y}^{\ast}M$ to $S_{x}^{\ast}M$. Concatenating these isotopies yields a non-negative Legendrian isotopy from $S^{\ast}_{x}M$ to itself. By [11, Remark 1.1] this proves the claim. ∎
3.28 Remark.
There is a different way to construct a non-negative Legendrian isotopy from one fiber of $S^{\ast}M$ to itself in the case where $M$ is the Cauchy surface of a globally hyperbolic strongly refocusing spacetime $(X,g)$. Let $p,q\in X$ have the same sky. By [14, Proposition 1.1], [1, Proposition 5.8] there is a non-negative Legendrian isotopy between skies of any causally related points. Since $p$ and $q$ are causally related with the same sky, there is a non-negative Legendrian isotopy from $\mathrm{sky}_{p}$ to itself. Fixing some Cauchy surface $M\subset X$ with $p\in M$ and identifying $\mathcal{N}$ with $S^{\ast}M$ we get a non-negative Legendrian isotopy from a fiber of $S^{\ast}M$ to itself.
The question about the cohomology of the universal cover of the Cauchy surface was an open question in [12]. The affirmative answer to it was known to Chernov and Nemirovski but not yet published.
@@ 4 Observer-Refocusing Spacetimes
Since the result is entirely local, we can assume $M=\mathbb{R}^{n}$ and $x=0$. Without loss of generality, assume that $x_{n}\to x$ and that $x_{n}\neq x$ for all $n$, else take a subsequence. Further, without loss of generality, assume that $\lim_{n\to\infty}\frac{x_{n}}{||x_{n}||}=v$ for some unit vector $v$ and that $||x_{n+1}||<\frac{1}{2^{n}}||x_{n}||$ for all $n$, else we again take a subsequence. Then the sequence $(w_{n})_{n\in\mathbb{N}}$ defined through $w_{n}\coloneqq\frac{x_{n}-x_{n+1}}{||x_{n}-x_{n+1}||}$ converges to $v$. We define the curve $\gamma$ as follows. We connect $x_{n+1}$ to $x_{n}$ by a $\mathcal{C}^{1}$ curve $\gamma_{n}:[0,a_{n}]$, parametrized by arc-length, satisfying $\gamma_{n}(0)=x_{n+1}$, $\gamma_{n}(a_{n})=x_{n}$, $\gamma_{n}^{\prime}(0)=w_{n+1}$ and $\gamma_{n}^{\prime}(a_{n})=w_{n}$. Since the sequence $(w_{n})_{n\in\mathbb{N}}$ converges to $v$, the interpolations can be done in a way (for example by interpolating with cubic splines) such that that $\lim_{n\to\infty}\sup_{t\in[0.a_{n}]}||\gamma_{n}^{\prime}(t)-v||=0$ and that $\sum_{n\in\mathbb{N}}a_{n}<\infty$. We then define $\gamma$ as the infinite concatenation
@@
Let $(X,g)$ be a Lorentzian manifold and let $h:[0,1]\to T_{p}X$ be a curve such that $h(t)$ is lightlike for each $t\in[0,1]$. Then $h^{\prime}(0)$ is not timelike. Hence $d\exp_{0}(h^{\prime}(0))=\frac{d}{dt}|_{t=0}\exp(h(t))$ is not timelike.
@@
To show that the Cauchy surface of $(X,g)$ has finite fundamental group, we proceed as follows. Let $\rho:(\tilde{X},\tilde{g})\to(X,g)$ be the semi-Riemannian universal cover of $X$. Let $\tilde{p}$ be any point in $\rho^{-1}(p)$ and let $\tilde{L}$ be the lightsphere at $p$ given by $\tilde{L}=(d\rho)^{-1}(L)$. Note that the set $\rho^{-1}(\{q_{i}\}_{i\in I}\})$ is countable. Let $\tilde{v}\in\tilde{L}$ and $v=d\rho(\tilde{v})\in L$. Then there is some $t\in(0,T]$ and $i\in I$ with $\alpha_{v}(t)=q_{i}$. Lifting $\alpha_{v}$ to $\tilde{X}$, we get that $\alpha_{\tilde{v}}(t)\in\rho^{-1}(q_{i})\in\rho^{-1}(\{q_{i}\}_{i\in I}\})$. Note that $\rho^{-1}(\{q_{i}\}_{i\in I}\})$ is countable. Through application of the first part of this theorem, we can conclude that $\tilde{X}$ has compact Cauchy surface as well. Now pick a Cauchy surface $S\subset X$ and let $\tilde{S}=\rho^{-1}(S)$. Then $\tilde{S}$ is a Cauchy surface of $\tilde{X}$ (see the proof of [15, Theorem 11.5]) and hence compact. Further, $\tilde{S}$ is simply connected (as $\tilde{X}$ is simply connected and diffeomorphic to $\tilde{S}\times\mathbb{R}$), and $\rho:\tilde{S}\to S$ is a covering map. Pick any $x\in\tilde{S}$, then
@@
It is unknown whether there are $Z^{x}$ manifolds where the time-of-return of unit-speed geodesics is not uniformly bounded [9, Question 7.71]. Similarly it is unknown whether there are any globally hyperbolic observer-refocusing spacetimes where the timelike curve cannot be taken to be compact.
@@ 5 The Analytic Case
Let $(X,g)$ be a globally hyperbolic spacetime with $g$ being an analytic metric. Then there exists an analytic, surjective temporal function $\mathcal{T}:X\to\mathbb{R}$ such that every level set of $\mathcal{T}$ is a spacelike Cauchy hypersurface.
@@ 6 Analogous conjecture in the Contact case
The results of Chernov and Nemirovski [13], the results of Frauenfelder, Labrousse and Schlenk [16, Theorem 1.11 and 1.13], and our results about Riemannian $Z^{(x,y)}$ manifolds and observer-refocusing spacetimes in this paper, suggest the following conjectures.
@@
Let $M$ be manifold of dimension at least $2$, let $(S^{\ast}M,\xi)$ be the spherical cotangent bundle on $M$ equipped with the standard contact structure. Let $\alpha$ be any contact form on $(S^{\ast}M,\xi)$, i.e. $\xi=\ker(\alpha)$, let $\varphi_{\alpha}^{t}$ be the Reeb flow of $\alpha$ and let $x,y\in M$. Assume that for all $v\in S^{\ast}_{x}M$ there is some $t>0$ with $\varphi_{\alpha}^{t}(v)\in S^{\ast}_{y}M$. Then $M$ is compact, the fundamental group of $M$ is finite, and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
@@
Let $M$, $\xi$, $\alpha$, $\varphi_{\alpha}^{t}$ be as in Conjecture 6.49. Let $x,y\in M$. Assume there is some $T>0$ such that for all $v\in S^{\ast}_{x}M$ there is some $t\in(0,T]$ with $\varphi_{\alpha}^{t}(v)\in S^{\ast}_{y}M$. Then $M$ is compact, the fundamental group of $M$ is finite and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
@@
This contact-geometric version generalizes the Riemannian $Z^{(x,y)}$ version as follows: Let $(M,g)$ be a Riemannian manifold, let $\psi:TM\to T^{\ast}M$ be the isomorphism defined through $v\mapsto g(v,\cdot)$, then $\psi$ induces a bundle-isomorphism (which we will also denote by $\psi$) between the unit-tangent bundle $U_{g}M$ and the spherical cotangent bundle $S^{\ast}M$. We can define a $1$-form $\alpha$ on $S^{\ast}M$ via $\alpha_{v}(w)=g(\psi^{-1}(v),d\pi(w))$ where $\pi:S^{\ast}M\to M$ is the natural projection. The Reeb flow of $\alpha$ on $S^{\ast}M$ corresponds to the co-geodesic flow of $g$ on $S^{\ast}M$. Conjecture 6.49 generalizes the conjecture that $Z^{x}$ manifolds of dimension at least $2$ have finite fundamental group. We have proven the compactness and the fundamental group aspect of Conjecture 6.50 whenever the contact form $\alpha$ comes from a Riemannian metric. The conjecture states that this is really a fact about Reeb flows on $(S^{\ast}M,\xi)$ and not just about geodesics on $M$. One may also ask other questions, for example if the set $R_{x}^{\alpha}\coloneqq\{t>0\mid\exists v\in S^{\ast}_{x}M\text{ such that }\varphi_{\alpha}^{t}(x)\in S^{\ast}_{x}M\}$ is discrete for all contact forms $\alpha$ on $(S^{\ast}M,\xi)$. If it is always discrete, then Conjecture 6.49 holds for analytic $1$-forms $\alpha$ by the analyticity of the Reeb flow and by the analyticity of $\pi:S^{\ast}M\to M$ because arguments as in Theorem 5.47 can reduce the problem to the case already solved by Frauenfelder, Labrousse and Schlenk.
6.51 Remark.
For a given globally hyperbolic spacetime $(X,g)$ there are many different splittings of $X=M\times\mathbb{R}$ like in Proposition 3.26, each giving rise to positive Legendrian isotopies $\phi_{t}$ of $S^{\ast}M$. In particular, if $\gamma$ is any timelike curve in $X$, it is always possible to split $X$ as $M\times\mathbb{R}$ such that in these coordinates $\gamma(t)=(y,t)$, i.e. such that $\gamma$ stays constant in the space-coordinate of the chosen coordinate system. Therefore, whenever we have a globally hyperbolic, observer-refocusing spacetime $X$ with Cauchy surface $M$, there will be $x,y\in M$ and a positive Legendrian isotopy $\phi_{t}$ on $S^{\ast}M$ such that for all $v\in S^{\ast}_{x}M$ there is some $t>0$ with $\phi_{t}(v)\in S_{y}^{\ast}M$.
The following two conjectures further generalize Conjecture 6.49 and Conjecture 6.50 by replacing Reeb flows with positive Legendrian isotopies.
6.52 Conjecture.
Let $M$ be manifold of dimension at least $2$, let $(S^{\ast}M,\xi)$ be the spherical cotangent bundle on $M$ equipped with the standard contact structure. Let $\phi_{t}$ be any positive Legendrian isotopy of $(S^{\ast}M,\xi)$ with $\phi_{0}=\mathrm{id}$ and let $x,y\in M$. If for all $v\in S^{\ast}_{x}M$ there is some $t>0$ with $\phi_{t}(v)\in S^{\ast}_{y}M$. Then $M$ is compact, the fundamental group of $M$ is finite, and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
6.53 Conjecture.
Let $M$, $\xi$, $\alpha$ and $\phi_{t}$ be as in Conjecture 6.52. Let $x,y\in M$. Assume there is some $T>0$ such that for all $v\in S^{\ast}_{x}M$ there is some $t\in(0,T]$ with $\varphi_{t}(v)\in S^{\ast}_{y}M$. Then $M$ is compact, the fundamental group of $M$ is finite and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
We have proven the compactness and fundamental group aspect of Conjecture 6.53 whenever the positive Legendrian isotopy comes from the construction seen in Proposition 3.26. The conjecture states that this is really a fact about positive Legendrian isotopies of $(S^{\ast}M,\xi)$ and not just about the null-geodesic flow in spacetimes which have $M$ as a Cauchy surface.